OpenStudy (anonymous):

suppose a triangle has two sides of length 32 and 35, and that the angle between these two sides is 120 degrees. what is the length of the third side of the triangle?

OpenStudy (jdoe0001):

have you covered the Law of Cosines yet?

OpenStudy (anonymous):

yeah but im having trouble understanding it

OpenStudy (jdoe0001):

 \(\bf \textit{Law of Cosines}\\ \quad \\ c^2 = a^2+b^2-(2ab)cos(c)\\ c = \sqrt{a^2+b^2-(2ab)cos(c)}\qquad {\color{red}{ a}}=32\qquad {\color{blue}{ b}}=35\qquad \measuredangle {\color{green}{ C}}=120^o\\ \quad \\ c = \sqrt{{\color{red}{ 32}}^2+{\color{blue}{ 35}}^2-(2\cdot {\color{red}{ 32}}\cdot {\color{blue}{ 35}})cos({\color{green}{ 120^o}})}\)

OpenStudy (anonymous):

flutter math